OpenStudy (anonymous):

Factorizing x^4+3x^2+2?

OpenStudy (jdoe0001):

\(\bf x^4+3x^2+2\implies ({\color{blue}{ x^2}})^2+3{\color{blue}{ x^2}}+2\qquad \qquad {\color{blue}{ u=x^2}}\quad thus \\ \quad \\ \begin{array}{cccllll} {\color{blue}{ u}}^2&+3{\color{blue}{ u}}&+2\\ &\uparrow &\uparrow \\ &2+1&2\cdot 1 \end{array}\)

OpenStudy (anonymous):

Thank you so much

OpenStudy (jdoe0001):

yw